OpenStudy (anonymous):

Given that 3*10^3 + 2*10^1 + 4*10^x +5*10^y = 3024.05, find x and y @arabpride

OpenStudy (arabpride):

wow so complex .. >.> okayy im going to try at least.. maybe I'll learn somethin while im at it cx

OpenStudy (anonymous):

thnks umm u know how to solve in standard form??

OpenStudy (campbell_st):

well isn't simply a case os 3000 + 20 + 4 + 0.05 so 3 x 10^3 = 3000 2 x 10^1 = 20 4 x 10^x = 4 and 5 x 10^y = 0.05 should be straight forward

OpenStudy (campbell_st):

what power of 10 gives 4 units...? and what power of 10 gives 1/100

OpenStudy (anonymous):

wait how'd u got 4*10^x??

OpenStudy (campbell_st):

well what you need to do is match the left side with the right side... the left side is the sum of powers of 10... so scientific notation. the right side is standard form... so its a matching exercise

OpenStudy (campbell_st):

10^3 = 1000 so 3 x 10^3 = 3000 10^1 = 10 so 2 x 10^1 = 20 10^x = 1 4 x 10^x = 4 does this help

OpenStudy (anonymous):

3020 + 4*10^X + 5*10^y =3024.05

OpenStudy (campbell_st):

that's correct...

OpenStudy (anonymous):

4*10^X +5*10^Y = 3024.05 - 3020??

OpenStudy (anonymous):

4*10^X + 5*10^Y = 4.05??

OpenStudy (campbell_st):

you don't need to do that.... do you know about place values.... e.g. 325 = 3 x 100^2 + 2 times 10^1 + 5 x 10^0

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

x=1

OpenStudy (campbell_st):

oops 325 = 3 x 10^2 + 2 x 10^1 + 5 x 10^0

OpenStudy (campbell_st):

nope... have a look at the example.... what is the power of 10 for the units...?

OpenStudy (anonymous):

power 2, power 1 and power 0

OpenStudy (anonymous):

x=0??

OpenStudy (anonymous):

and y=0.5??

OpenStudy (campbell_st):

that's correct... here it is as a diagram of powers of 10  what do you think the values of x and y are..? look at the powers